OpenStudy (anonymous):

simplify x^2-3x-18 over x^2 -36

OpenStudy (anonymous):

\[\frac{x^2-3x-18}{x^2-36} \text{ = }\frac{(x-6) (x+3)}{(x-6) (x+6)}\text{ = }\frac{x+3}{x+6} \]